A block of mass m bogins at rest at the top of a r
Answer #1

The data given in the question is,

m=7.8kg,h=8.1m,d=5m,\mu=0.3,angle=36.87degree

Potential energy PE=mgh=7.8*9.8*8.1=619.2J

Frictional force =mgsin\Theta\muk =7.8*9.8*sin(36.87)*0.3=14N

work done by friction=force* dstance=14*5=70J

work done by friction= change in total energy

work done by friction=change in KE+change in PE

Remaining energy to be converted to KE is, difference between work done by frictionand PE

Remaining energy to be converted to KE=619.2-70= 549.2 J